OpenStudy (anonymous):

Write the equation of a line that passes through the points (2,4) (1,-3) in slope-intercept form

OpenStudy (anonymous):

use formula... m=(y1-y)/(x1-x) m=(4-(-3))/(2-1) m=7/1=7 then equation of the line: y-4=7(x-2) y=7x-14+4 y=7x-10

OpenStudy (anonymous):

Thanks,what about this one.. Write the equation of the line that passes through (1, 3) and (4, 4) in standard form.

OpenStudy (anonymous):

the same approach as above...

OpenStudy (anonymous):

m=(4-3)/(4-1) m=1/3 y-3=1/3 (x-1) y=1/3 x -1/3 +3 y=1/3 x +8/3 standard form - multiply both sides by 3 & write all on one side: x-3y+8=0

OpenStudy (anonymous):

You really are a lifesaver (: can i bother you for one last one? Part 1: Write the equation of the line that passes through the points (3, 4) and (-3, -8) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer.

OpenStudy (anonymous):

Part2: to get the final answer - equation of the line it is NOT matter which point is used. you can first one (see above) or second one the answer should be the same

OpenStudy (anonymous):

Part1: m=(4-(-8))/(3-(-3))=12/6 m=2 y-4=2 (x-3) y=2 x -6+4 y=2 x -2 this is in slope-intercept form

OpenStudy (anonymous):

YOU can try second point - should be the same answer: y-(-8)=2(x-(-3)) y=2x+6-8 y=2x-2 the same!